# Reducing Size Bias in Sampling Infectious Disease Spread on Networks

### v2

@@ Appendix A Pseudocode for sampling algorithms
We provide the logic of the code for extracting samples using the Random Walk (RW) (Algorithm 1) and Metropolis-Hastings Random Walk (MHRW) (Algorithm 2) algorithms; refer to Section 2.2 for the mathematical formulation of the algorithms. For both algorithms, to obtain a sample from a graph, we provide three inputs: i) a graph object, which contains the index or label of nodes, and the list of connections between nodes, ii) the number of samples, and iii) the sample size, which is the number of nodes required in a sample.
@@ Appendix B Statistical test results
We discuss the results of the normality test on the data for two disease metrics: proportion of infected nodes, and number of secondary infections. We perform the normality test with the hypotheses stated in Section 2.5 on the data of disease metrics from the UN, the RW samples, and the MHRW samples. First, we assess whether the data follows a Normal distribution or not, as it influences the choice of statistical tests for comparing the disease metrics among the RW, MHRW, and UN data. If the data is normally distributed then parametric tests will be used; otherwise, non-parametric tests will be used.
We use the one-sample Kolmogorov-Smirnov (KS) test with a significance level of $\alpha=0.05$ to assess the normality hypotheses (2.1) and (2.2). The KS test statistic is the measure of the largest absolute difference between the cumulative distribution of the data (UN, RW, MHRW) and the normal distribution, which ranges from 00 to $1$. A larger KS statistic indicates a greater deviation from the normal distribution. A KS statistics value closer to $1$ with a p-value $<0.05$, suggests that the data distribution is significantly different from the normal distribution, while a value closer to 00 with a p-value $>0.05$ implies that the data is significantly similar to a normal distribution. Below are the results for the two disease metrics across three networks (ER, SW, and SF) for the data from the UN, RW, and MHRW samples.
@@
    Proportion Infected Nodes: For all three networks, the KS statistic (Table 5) values are approximately $0.5$, with p-values $<0.05$, evidence to reject the $H_{0}$ in Hypothesis (2.1) for the UN, RW and MHRW data. The KS statistic is sufficiently large, leading us to conclude that the data distributions for the proportion of infected nodes significantly deviate from a Normal distribution.
@@
    Secondary Infections: For all three networks, the KS statistic (Table 5) values are approximately $0.5$, with p-values $<0.05$, evidence to reject the $H_{0}$ in Hypothesis(2.2) for the UN, RW and MHRW data. The KS statistic is sufficiently large, leading us to conclude that the data distributions for the number of secondary infections significantly deviate from a Normal distribution.
@@
Table 4: One-sample KS test statistics for the proportion of infected nodes (Hypothesis (2.1)).
@@
Table 5: One-sample KS test statistics for the number of secondary infections (Hypothesis (2.2)).
Since, data does not follow a Normal distribution across networks, for both disease metrics, a non-parametric statistical test is required to compare the RW and MHRW samples with the UN. We are using the Man-Whitney U-test in this work, results for the same are discussed in Section 3.
@@ Appendix C Disease metrics
We compare the sampling algorithms (RW and MHRW) based on the accuracy of disease metric estimates for the ER, SW and SF networks using the method described in Section 2.4. Specifically, we discuss the estimates of the time-to-infection for increasing $\beta$ values while keeping $\gamma=1$ in Section C.1. In Section C.2, we compare the sampling algorithms for three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time-to-infection for the ER and SW networks, with $\beta$ values from 00 to $1$ and $\gamma\in\{1/7,1/14\}$.
C.1 Time to get infection
In Figure 8, we compare the estimates of the time-to-infection from the RW and MHRW samples with the UN for increasing $\beta$ values and $\gamma=1$ for three networks (ER, SW, and SF). In Panel 1 (Figure 7(a), 7(b), 7(c)) results for samples with duplicate nodes and in Panel 2 (Figure 7(d),7(e),7(f)) results for samples without duplicate nodes are shown.
We observe that for ER networks and $\beta<0.2$, there is a large variation in the estimates of both the sampling algorithms in both panels. For $\beta\geq 0.2$, the MHRW sample estimates align well with the UN, while the RW sample estimates slightly underestimate the time-to-infection, if duplicate nodes are retained in the samples (Figure 7(a)). When duplicate nodes are removed, the MHRW sample estimates slightly underestimate the time-to-infection (Figure 7(d)), although the behaviour of the RW sample estimates remains consistent. For SW networks, a behaviour similar to that of the ER networks is observed. However, the large variation in the estimates for both sampling algorithms persists up to $\beta=0.5$.
Surprisingly, for SF networks, as shown in Figures 7(c) and 7(f), the behaviour of both algorithms remains consistent across both panels, i.e., estimates from both the algorithms align well with the UN and have small variation band.
Overall, both sampling algorithms are performing well across network types, particularly for the Scale-free (SF) networks, which is not the case for the other two disease metrics (the proportion of infected nodes and the number of secondary infections).
Panel 1: Samples with duplicate nodes.
(a) Erdős-Rényi network
(b) Small-world network
(c) Scale-free network
Panel 2: Samples without duplicate nodes.
(d) Erdős-Rényi network
(e) Small-world network
(f) Scale-free network
Figure 8: Comparison of average time to get infected between samples generated using RW (dashed red line) and MHRW (dotted blue line), relative to the UN. (1) Duplicate nodes in the samples are retained, and (2) duplicate nodes in the samples are removed. The light red and blue bands indicate the range of one standard deviation. The UN estimates are based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
C.2 Sensitivity analysis - Recovery rate
In Section 3, we discussed the results with a varying transmission rate $\beta$ and a fixed $\gamma=1$; here we investigate the impact of varying $\gamma$ along with $\beta$ on the accuracy of disease estimates from sampling algorithms for the ER and SW networks. We compare the disease metric estimates from both sampling algorithms (RW and MHRW) with the UN, using three values of recovery rate $\gamma$ that are $1$, $1/7$, and $1/14$. We also discuss the results based on two scenarios: with duplicate nodes in the samples, and without duplicate nodes in the samples.
C.2.1 Samples with duplicate nodes
In Figure 9, we compare the average ratio of disease metric estimates between the RW and MHRW samples (with duplicate nodes) with the UN for three disease metrics: i) proportion of infected nodes, ii) number of secondary infections, and iii) time to get infected in days. The recovery rate ($\gamma$) values are $1,1/7$, and $1/14$ and transmission rate $\beta$ increases from 00 to $1$. The purple line plots with circle markers is for $\gamma=1$, the brown line plots with triangle markers is for $\gamma=1/7$, and the green line plots with square markers is for $\gamma=1/14$. For a network and a disease metric, there are two plots, one for the RW sample estimates and another for the MHRW sample estimates.
@@ C.2.1 Samples with duplicate nodes
For the ER network, as shown in Figures 8(e) and 8(f), the sample estimates for the number of secondary infections deviate far from the UN as the recovery rate decreases from $1$ to $1/14$, for both sampling algorithms. The MHRW sample estimates are closer to the UN compared to the RW sample estimates irrespective of the recovery rate values. We observe similar behaviour for the SW networks, as shown in Figures 8(g) and 8(h).
For the estimates of the time-to-infection in the ER network, as shown in Figures 8(i) and 8(j), the sample estimates deviate away from the UN for both sampling algorithms as recovery rate decreases. The overestimation increases for the MHRW samples and the underestimation increases for the RW sample estimates. For the SW networks, as shown in Figures 8(k) and 8(l), underestimation increases for the RW sample estimates. Whereas the MHRW sample estimates align well with the UN with the decrease in recovery rate from $1$ to $1/14$ for $\beta<0.4$.
@@
Figure 9: The average value of the ratio of three disease metrics estimates for ER and SW networks using RW and MHRW samples (with duplicate nodes) relative to the UN; for varying recovery rate $\gamma\in\{1,1/7,1/14\}$ and transmission rate $\beta$ from 00 to $1$. The UN estimates are based on SIR model simulations for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
C.2.2 Samples without duplicate nodes
In Figure 10, we compare the average ratio of disease metric estimates between the RW and MHRW samples (without duplicate nodes) with the UN for three disease metrics: i) proportion of infected nodes, ii) number of secondary infections, and iii) time to get infected in days. The recovery rate ($\gamma$) values are $1,1/7$, and $1/14$, and transmission rate $\beta$ increases from 00 to $1$. The purple line plots with circle markers is for $\gamma=1$, the brown line plots with triangle markers is for $\gamma=1/7$, and the green line plots with square markers is for $\gamma=1/14$. For a network and a disease metric, there are two plots, one for the RW sample estimates and another for the MHRW sample estimates.
@@ C.2.2 Samples without duplicate nodes
Figure 10: The ratio of the average value of three disease metrics estimates for ER and SW networks using RW and MHRW samples (without duplicate nodes) relative to the UN; for varying recovery rate $\gamma\in\{1,1/7,1/14\}$ and transmission rate $\beta$ from 00 to $1$.The UN estimates are based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (parameters in Table 1) generated with NetworkX (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
@@
For the estimates of the time-to-infection in the ER network, as shown in Figures 9(i) and 9(j), we observe that sample estimates deviate away from the UN for both sampling algorithms as the recovery rate decreases. The underestimation increases for the RW sample estimates and overestimation increases for the MHRW sample estimates. For the SW networks, as shown in Figures 9(k) and 9(l), underestimation increases for the RW sample estimates. Whereas the MHRW sample estimates align well with the UN with the decrease in recovery rate from $1$ to $1/14$ for $\beta<0.4$.
Overall, from comparing the sample estimates for three $\gamma$ values ($1$, $1/7$, and $1/14$), we observe that the estimates for both sampling algorithms are very similar when $\gamma=1/7$ and $\gamma=1/14$. However, with $\gamma=1$, there is a noticeable deviation in the sample estimates. This indicates that the sample estimates are sensitive to the recovery rate, as well as to the type of disease metric and network type.

### v3

@@ Appendix A Pseudocode for Sampling Algorithms
We provide the logic of the code for extracting samples using the Random Walk (RW) (Algorithm 1) and Metropolis-Hastings Random Walk (MHRW) (Algorithm 2) algorithms; refer to Section 2.2 for the mathematical formulation of the algorithms. For both algorithms, to obtain a sample from a graph, we provide three inputs: i) a graph object, which contains the index or label of nodes and the list of connections between nodes; ii) the number of samples, and iii) the sample size, which is the number of nodes required in a sample.
@@ Appendix B Statistical test results
We discuss the results of the normality test on the data for two disease metrics: proportion of infected nodes and number of secondary infections. We perform the normality test with the hypotheses stated in Section 2.5 on the data of disease metrics from the UN, the RW samples, and the MHRW samples. First, we assess whether the data follows a Normal distribution or not, as it influences the choice of statistical tests for comparing the disease metrics among the RW, MHRW, and UN data. If the data is normally distributed, then parametric tests will be used; otherwise, non-parametric tests will be used.
We use the one-sample Kolmogorov-Smirnov (KS) test with a significance level of $\alpha=0.05$ to assess the normality hypotheses 2.1 and 2.2. The KS test statistic is the measure of the largest absolute difference between the cumulative distribution of the data (UN, RW, MHRW) and the normal distribution, which ranges from 00 to $1$. A larger KS statistic indicates a greater deviation from the normal distribution. A KS statistics value closer to $1$ with a p-value $<0.05$ suggests that the data distribution is significantly different from the normal distribution, while a value closer to 00 with a p-value $>0.05$ implies that the data is significantly similar to a normal distribution. Below are the results for the two disease metrics across three networks (ER, SW, and SF) for the data from the UN, RW, and MHRW samples.
@@
    Proportion Infected Nodes: For all three networks, the KS statistic (Table 6) values are approximately $0.5$, with p-values $<0.05$, evidence to reject the $H_{0}$ in Hypothesis 2.1 for the UN, RW and MHRW data. The KS statistic is sufficiently large, leading us to conclude that the data distributions for the proportion of infected nodes significantly deviate from a Normal distribution.
@@
    Secondary Infections: For all three networks, the KS statistic (Table 6) values are approximately $0.5$, with p-values $<0.05$, evidence to reject the $H_{0}$ in Hypothesis 2.2 for the UN, RW and MHRW data. The KS statistic is sufficiently large, leading us to conclude that the data distributions for the number of secondary infections significantly deviate from a Normal distribution.
  3. 3.
    Time to Infections: The KS statistic (Table 7) values are approximately $0.7$ for ER, $0.6$ for SW, and $0.5$ for SF, with p-values $<0.05$, evidence to reject the $H_{0}$ in Hypothesis 2.3 for the UN, RW and MHRW data. The KS statistic is sufficiently large, leading us to conclude that the data distributions for the number of secondary infections significantly deviate from a Normal distribution.
@@
Table 5: One-sample KS test statistics for the proportion of infected nodes (Hypothesis 2.1).
@@
Table 6: One-sample KS test statistics for the number of secondary infections (Hypothesis 2.2).
Network UN RW MHRW
ER 0.729 0.721 0.726
SW 0.659 0.663 0.662
SF 0.500 0.500 0.500
Table 7: One-sample KS test statistics for the number of secondary infections (Hypothesis 2.3).
Since, data does not follow a Normal distribution across networks, for all three disease metrics, a non-parametric statistical test is required to compare the RW and MHRW samples with the UN. We are using the Man-Whitney U-test in this work; results for the same are discussed in Section 3.
@@ Appendix C Disease metrics
We compare the sampling algorithms (RW and MHRW) based on the accuracy of disease metric estimates for the ER, SW and SF networks using the method described in Section 2.4. Specifically, in Section C.1, we compare the sampling algorithms for three disease metrics: 1) proportion of infected nodes, 2) average number of secondary infections, and 3) time-to-infection for the ER and SW networks, with $\beta$ values from 00 to $1$ and $\gamma\in\{1/7,1/14\}$.
C.1 Sensitivity analysis - Recovery rate
[…]
C.1.1 Samples with duplicate nodes
In Figure 9, we compare the average ratio of disease metric estimates between the RW and MHRW samples (with duplicate nodes) with the UN for three disease metrics: i) proportion of infected nodes, ii) number of secondary infections, and iii) time to get infected in days. The recovery rate ($\gamma$) values are $1,1/7$, and $1/14$ and transmission rate $\beta$ increases from 00 to $1$. The purple line plots with circle markers are for $\gamma=1$, the brown line plots with triangle markers are for $\gamma=1/7$, and the green line plots with square markers are for $\gamma=1/14$. For a network and a disease metric, there are two plots, one for the RW sample estimates and another for the MHRW sample estimates.
@@ C.1.1 Samples with duplicate nodes
For the ER network, as shown in Figures 8(e) and 8(f), the sample estimates for the number of secondary infections deviate far from the UN as the recovery rate decreases from $1$ to $1/14$, for both sampling algorithms. The MHRW sample estimates are closer to the UN compared to the RW sample estimates, irrespective of the recovery rate values. We observe similar behaviour for the SW networks, as shown in Figures 8(g) and 8(h).
For the estimates of the time-to-infection in the ER network, as shown in Figures 8(i) and 8(j), the sample estimates deviate away from the UN for both sampling algorithms as recovery rate decreases. The overestimation increases for the MHRW samples, and the underestimation increases for the RW sample estimates. For the SW networks, as shown in Figures 8(k) and 8(l), underestimation increases for the RW sample estimates. At the same time, the MHRW sample estimates align well with the UN, showing a decrease in the recovery rate from $1$ to $1/14$ for $\beta<0.4$.
@@
Figure 9: The average value of the ratio of three disease metrics estimates for ER and SW networks using RW and MHRW samples (with duplicate nodes) relative to the UN; for varying recovery rate $\gamma\in\{1,1/7,1/14\}$ and transmission rate $\beta$ from 00 to $1$. The UN estimates are based on SIR model simulations for $10,000$ networks of each type (the parameters are found in Table 1) generated with the NetworkX library (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
C.1.2 Samples without duplicate nodes
In Figure 10, we compare the average ratio of disease metric estimates between the RW and MHRW samples (without duplicate nodes) with the UN for three disease metrics: i) proportion of infected nodes, ii) number of secondary infections, and iii) time to get infected in days. The recovery rate ($\gamma$) values are $1,1/7$, and $1/14$, and transmission rate $\beta$ increases from 00 to $1$. The purple line plots with circle markers are for $\gamma=1$, the brown line plots with triangle markers are for $\gamma=1/7$, and the green line plots with square markers are for $\gamma=1/14$. For a network and a disease metric, there are two plots, one for the RW sample estimates and another for the MHRW sample estimates.
@@ C.1.2 Samples without duplicate nodes
Figure 10: The ratio of the average value of three disease metrics estimates for ER and SW networks using RW and MHRW samples (without duplicate nodes) relative to the UN; for varying recovery rate $\gamma\in\{1,1/7,1/14\}$ and transmission rate $\beta$ from 00 to $1$.The UN estimates are based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (the parameters are found in Table 1) generated with the NetworkX library (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
@@
For the estimates of the time-to-infection in the ER network, as shown in Figures 9(i) and 9(j), we observe that sample estimates deviate away from the UN for both sampling algorithms as the recovery rate decreases. The underestimation increases for the RW sample estimates, and the overestimation increases for the MHRW sample estimates. For the SW networks, as shown in Figures 9(k) and 9(l), underestimation increases for the RW sample estimates. At the same time, the MHRW sample estimates align well with the UN, showing a decrease in the recovery rate from $1$ to $1/14$ for $\beta<0.4$.
Overall, from comparing the sample estimates for three $\gamma$ values ($1$, $1/7$, and $1/14$), we observe that the estimates for both sampling algorithms are very similar when $\gamma=1/7$ and $\gamma=1/14$. However, with $\gamma=1$, there is a noticeable deviation in the sample estimates. This suggests that the sample estimates are sensitive to both the recovery rate and the type of disease metric, as well as the network type.